# Inverse design of multiresonance filters via quasi-normal mode theory

### v1

@@ I Introduction
To address these issues, we recently derived simple and general analytical rules to design standard filters [24] based on our earlier introduced quasi-normal mode theory (QNMT) for the scattering matrix $S$ [25]. In particular, we showed that the sharp (of high quality factor $Q$) 2-port-system resonances, apart from having eigenfrequencies matching the poles of the desired SF, must also have eigenfields coupling to the two ports with specific unitary ratios. In addition, an appropriate amount of background transmission must be enforced to generate the desired overall transmission spectrum away from the resonances. Essentially, the success of the method relies on the fact that, through QNMT, the rapidly oscillating part of the spectrum is designed by setting objectives on the resonances rather than the transmission itself, so the remaining slowly varying background term can be optimized more easily. A review of our QNMT [25] and filter-design criteria [24] is presented in Sec. II. Still, it is not immediately evident how to incorporate these analytical criteria into a numerically tractable formulation of an optimization problem. Previous TopOpt efforts to design cavity resonances typically maximized the power exerted by a dipole source, or equivalently the local density of states (LDOS), at a single real excitation frequency [27, 42, 43]. However, such an approach creates only one resonance and only at a frequency with real part equal to the excitation frequency, without being able to control the QNM’s decay rate (imaginary part of the pole) nor its field profile (thus its coupling to the ports). Some other works optimized for resonances by instead targeting a desired transmission spectrum, but were limited to only a single resonance [44] or resonances at several spectrally well-separated frequencies (non-overlapping peaks) [45]. Our previous work where the analytical criteria were derived [24] utilized an eigenvalue solver to directly optimize the $N$ high-$Q$ resonances for an $N$th-order filter. However, for a random initial structure, it is not clear which $N$ of the initial (likely low-$Q$) eigenvalues are to be optimized. Similarly, it is not always easy to guess a reasonable initial structure with $N$ high-$Q$ resonances in the vicinity of the desired targets (in [24], we used intuition from circuit theory to guess a reasonable topology and its initial geometrical parameters). Furthermore, even if $N$ high-$Q$ poles are identified, as the optimization progresses and they shift, one needs to track them smoothly (differentiably) and this is not always practical, especially when two poles cross in the complex plane. Finally, compared to scattering solvers, eigensolvers tend to be slow, since they rely on solving a sequence of many linear systems [46].
@@ II Quasi-normal mode theory and inverse-design criteria
The QNMT developed in [25] considers a multiport scattering system, whose outgoing waves are related to the incoming excitations at frequency $\omega$ by a scattering matrix $S(\omega)$, and provides a simple analytical model for calculating $S(\omega)$ with very high accuracy using only knowledge about the system resonant modes. The $N$ system high-$Q$ resonances (QNMs) have complex frequencies $\omega_{n}$, which appear as poles of $S$, and complex coupling coefficient to each port $p$ equal to $D_{pn}$, which can be computed as an overlap surface integral [50] between the $n$-QNM field and the $p$-port mode at the boundary of the scatterer, as explained in detail in [25]. Conveniently, the ratios $D_{pn}/D_{qn}$ do not depend on the normalization of the QNMs ([25], Appendix D). QNMT factors these $N$ high-$Q$ resonances into a pole-expansion sum, indicated as $\bar{S}(\omega)$ below. Additionally, low-$Q$ resonances (with poles $\omega_{n}^{C}$ and couplings $\sigma_{n}^{C}$) can be factored into an effective-background matrix $C(\omega)$, which varies slowly with frequency (approximately constant in most practical design cases). For the case of a 2-port system, the ratios $\sigma_{n}=D_{2n}/D_{1n}$ were defined and a schematic is shown in Fig. 1. The key result from [25] is that, for a lossless system, by enforcing energy conservation (unitary $S$) and maximizing reciprocity (fine-tuning $\sigma_{n}$ to also best symmetrize $S$ — exactly symmetric when $C$ constant), the $2\times 2$ scattering matrix $S(\omega)$ has a simple, approximate, but very accurate expression that only depends on $\{\omega_{n},\sigma_{n}\}$ and $C$:
@@
(Here, $\bar{S}$ is defined as the indicated sum; note that a bar $\bar{\square}$ does not denote complex conjugation of $\square$, for which we use an asterisk $\square^{*}$.) The background $C$ is itself unitary symmetric and could conceptually be computed from the remaining (presumably low-$Q$) system poles: $C(\omega)=-\bar{S}_{\{\omega_{n}^{C},\sigma_{n}^{C}\}}(\omega)$. When no separate background is needed (all modes are considered as high-$Q$), then $C=-I$. Moderate losses can also be incorporated perturbatively [25] by adjusting the poles only in the denominator of the $\bar{S}$ expansion in Eq. (1).
@@
For an arbitrary desired frequency response $\tilde{S}(\omega)$, one could determine the target QNMT parameters by fitting $\tilde{S}(\omega)$ with the QNMT model of Eq. (1). Because $\tilde{S}(|\omega/\omega_{n}-1|\gg 1)\rightarrow\tilde{C}$, one can usually infer $|\tilde{r}|$ by the desired response’s asymptotic amplitude away from the resonant features, at the edges of the frequency range of interest. The procedure for this fit is detailed in Appendix A.
For particular spectral responses, however, that correspond to standard $N$th-order filters (Butterworth, Chebyshev, inverse Chebyshev, and elliptic), described by rational functions of $i\omega$ with real coefficients [26], [24] has shown that the target QNMT parameters conveniently take specific analytical values: (i) The $N$ high-$Q$ complex poles $\tilde{\omega}_{n}$ are immediately available from the denominator of the desired-filter rational function [26], (ii) their coupling ratios $\tilde{\sigma}_{n}$ must have unity magnitude and alternate signs,
@@
$\tilde{r}=e^{-i\varphi}i^{N\mp 1}|\tilde{r}|,$ (3)
with the $\mp$ signs corresponding to bandpass and bandstop filters respectively, and $|\tilde{r}|$ is 1 for an odd-order bandpass transmission SF (asymptotes to full reflection in stopband), 0 for odd-order bandstop, large (close to 1) for even-order elliptic or inverse-Chebyshev bandpass, and small (close to 0) for even-order elliptic or inverse-Chebyshev bandstop. As explained earlier, for practical design problems where the RWA applies, $\varphi$ can be any phase and is thus a design degree of freedom, or alternatively a hyperparameter that may be chosen to generate different solutions.
@@
$\begin{gathered}\{\omega_{n},\sigma_{n}\}\rightarrow\{\tilde{\omega}_{n},\tilde{\sigma}_{n}\}\\ \text{and }C_{11}^{*}(\omega)C_{21}(\omega)\rightarrow\tilde{r}^{*}\tilde{t}\,.\end{gathered}$ (4)
@@ III.2 Enforcing background matrix
[…]
@@ IV.1 Physical model of metasurface scattering
$[-\nabla\cdot\frac{1}{\varepsilon(x,y)}\nabla-k_{o}^{2}\mu]H_{z}=J$ (8)
@@
[…]
@@ IV.2 Density-based topology optimization
Figure 3: Designed photonic 3rd-order elliptic bandpass filter: (a) The structure is shown on the top left, with the design region boxed inside the dashed orange line. Its transmission spectrum $|S_{21}(f)|^{2}$ is plotted as the orange curve, with an enlargement of the passband on the top right. Its background transmission $|C_{21}(f)|^{2}$ is below $-70$dB (and hence omitted from the plot). Therefore, the response of our design matches perfectly the target standard filter (black curve). The traditional naive-optimization method performs very poorly in comparison (green curve). (b) Target and designed poles and their coupling ratios $\{\omega_{n},\sigma_{n}\}$.
@@
$-r_{\text{f}}^{2}\nabla^{2}\rho_{\text{f}}+\rho_{\text{f}}=\rho,\quad\text{with}\;\vec{n}\cdot\nabla\rho_{\text{f}}=0\;\text{on}\;{\partial\Omega}$ (9)
@@
The specifications of all transmission filters designed below are: bandpass type with $1\%$ pass bandwidth around $\lambda=1$, $0.25$dB transmission ripple allowed in the passband, and at least $25$dB transmission suppression in the stopband. The two dielectric materials considered for the structure are $n_{\text{low}}=1$ (air) and $n_{\text{high}}=3.4$ (silicon). The incidence and transmission air regions are $0.5\lambda$ thick, which is also the thickness of the PMLs, with absorption large enough to reduce reflections to $10^{-12}$ at the left and right boundaries. To make the optimization of this bandpass filter easier, the initial structure was always chosen to be a quarter-wave Bragg stack at $\lambda=1$, namely alternating layers of $0.25$-thick air and $0.25/n_{\text{high}}$-thick silicon, so that the background transmission is close to its desired small value. To obtain reasonable computational accuracy, we chose a mesh density of about 6000 triangular mesh elements per $\lambda^{2}$ (so the average side of the triangular elements is roughly $0.02\lambda$, or $0.07\lambda/n_{\text{high}}$ in the material).
@@
As a first example, we designed a third-order ($N=3$) elliptic filter. For an odd-order filter, $\tilde{t}=0$, so the $C$ constraint in Eqs. (4) and (6) becomes $C_{21}(\omega_{m})\rightarrow 0$. The design region is chosen as $3\lambda$ long and $0.25\lambda$ high ($a=0.5\lambda$). The background transmission was constrained at $M=21$ frequency points, equally spaced within $[0.85,1.15]/\lambda$. After trying multiple values for $\alpha$ and $\varphi$, we obtained the best optimization result for $\alpha=0.02$ and $\varphi=\pi/2$, namely the coupling ratios are $\sigma_{n}=\{i,-i,i\}$ from Eq. (2). The relative errors of the actual poles and coupling ratios, compared with the targets, are within $10^{-5}$. Moreover, $C_{21}$ is less than $-70$dB, effectively zero, for a free-spectral range of $30\%$. Therefore, as shown in Fig. 3, we had an exceptionally good match between the actual transmission and the target spectrum.
@@
Figure 4: Designed photonic 4th-order elliptic bandpass filter: (a) The structure is shown on the top left. Its transmission spectrum $|S_{21}(f)|^{2}$ is plotted in orange, with an enlargement of the passband on the inset. Its background transmission $|C_{21}(f)|^{2}$ is steady around the specified $-25$dB. Again, our design exhibits perfect match to the target standard filter (black curve). (b) Poles of the designed structure: the high-Q poles and their coupling ratios $\{\omega_{n},\sigma_{n}\}$ shown in the enlarged inset match their targets, while the low-Q modes collectively account for the desired background transmission.
@@
As a second example, we designed a fourth-order ($N=4$) elliptic filter, in which (even-$N$) case the $-25$dB stopband limit sets the background transmission target to the non-trivial value $\tilde{t}=10^{-25/20}=0.0562$. We chose $\varphi=-\pi/2$, so the coupling ratios are $\sigma_{n}=\{-i,i,-i,i\}$ from Eq. (2) and the target background reflection coefficient is $\tilde{r}=0.9984+\text{i}0$ from Eq. (3). As before, the design region is made $3\lambda$ long and $0.25\lambda$ high ($a=0.5\lambda$). For the background constraint, we used $M=92$ frequencies equally spaced within $[0.8,1.2]/\lambda$, with $\alpha=1/150/\sqrt{M}$. After optimization, again the actual poles and coupling ratios have reached their targets to within $10^{-5}$, while $|C_{21}|^{2}$ is almost exactly at $-25$dB for a $40\%$ FSR. As shown in Fig. 4, again we have an amazing match of the actual transmission to the target spectrum.
@@ IV.4 Comparison to naive direct design
[…]
@@ V Few-parameter optimization
Figure 5: Designed 1D-layered photonic 3rd-order Chebyshev bandpass filter: The dielectric-layer stack is shown on the left, where black is Si and grey is SiO2, so light is normally incident from the top air and transmits into the bottom silica substrate; the $28$ layer thicknesses (from top to bottom, alternating SiO2–Si) are $d_{k}/\lambda=(0.3528,$ $0.07358,$ $0.1787,$ $0.07361,$ $0.3449,$ $0.08524,$ $0.1795,$ $0.07385,$ $0.1793,$ $0.07383,$ $0.1794,$ $0.07391,$ $0.1804,$ $0.03658,$ $0.04277,$ $0.07453,$ $0.1794,$ $0.07382,$ $0.1792,$ $0.07380,$ $0.1793,$ $0.07385,$ $0.1797,$ $0.1212,$ $0.2876,$ $0.07501,$ $0.1854,$ 0.2154)0.2154)0.2154 ). The transmission spectrum $|S_{21}(f)|^{2}$ (orange) matches the target filter (black) for a wide frequency range around the passband (enlarged in the inset). The background transmission $|C_{21}(f)|^{2}$ is below $-58$dB for the entire range.
As an example, we considered such a stack of dielectric films, alternating between silicon (Si with $n_{\text{high}}=3.4$) and silicon dioxide (SiO2 with $n_{\text{low}}=1.4$), and sitting on a SiO2 substrate. We designed a third-order ($N=3$) Chebyshev filter (again bandpass with $1\%$ bandwidth and $0.25$dB passband ripple) for plane-wave transmission from the air region above the stack into the substrate (assumed to be semi-infinite). We chose $\varphi=0$, namely the coupling ratios are $\sigma_{n}=\{1,-1,1\}$ from Eq. (2). [Note that this will force the tails of the QNMs to be (anti-)symmetric, although the structure is asymmetric.] We started with a 29-layer quarter-wave Bragg stack, namely $d_{k}=0.25\lambda/n_{k}$ for $k=1...29$, where $n_{k}$ is the refractive index of each layer, so $d_{k=\text{odd}}=0.0735$ for Si and $d_{k=\text{even}}=0.179$ for SiO2. We bounded the individual layer thicknesses to $0.75/n_{k}$ and the total Si thickness of Eq. (13) to $d_{\rho,\text{max}}=1.5N\lambda/n_{\text{high}}$ (with multiplier $\gamma=10$), while the total thickness of Eq. (12) was left unconstrained. The initial Bragg stack already gives very low transmission, matching our desired background, and it turned out that we did not need to impose any background constraints ($M=0$).
The optimization led to a structure where the top Si layer was less than $0.01/\lambda$ thick. Therefore, we removed this layer and repeated the optimization with the remaining 28 thicknesses. The resulting transmission spectrum is shown in Fig. 5. Once again, an excellent match between the target and actual poles and passband transmission is observed. $C_{21}$ is less than $-53$dB, sufficiently close to zero, for a free-spectral range of over $40\%$.

### v2

@@ I Introduction
To address these issues, we recently derived simple and general analytical rules to design standard filters [6] based on our earlier introduced quasi-normal mode theory (QNMT) for the scattering matrix $S$ [5]. In particular, we showed that the sharp (of high quality factor $Q$) 2-port-system resonances, apart from having eigenfrequencies matching the poles of the desired SF, must also have eigenfields coupling to the two ports with specific unitary ratios. In addition, an appropriate amount of background transmission must be enforced to generate the desired overall transmission spectrum away from the resonances. Essentially, the success of the method relies on the fact that, through QNMT, the rapidly oscillating part of the spectrum is designed by setting objectives on the resonances rather than the transmission itself, so the remaining slowly varying background term can be optimized more easily. A review of our QNMT [5] and filter-design criteria [6] is presented in Sec. II. Still, it is not immediately evident how to incorporate these analytical criteria into a numerically tractable formulation of an optimization problem. Previous TopOpt efforts to design cavity resonances typically maximized the power exerted by a dipole source, or equivalently the local density of states (LDOS), at a single real excitation frequency [48, 76, 33]. However, such an approach creates only one resonance and only at a frequency with real part equal to the excitation frequency, without being able to control the QNM’s decay rate (imaginary part of the pole) nor its field profile (thus its coupling to the ports). Some other works optimized for resonances by instead targeting a desired transmission spectrum, but were limited to only a single resonance [19] or resonances at several spectrally well-separated frequencies (non-overlapping peaks) [31]. Our previous work where the analytical criteria were derived [6] utilized an eigenvalue solver to directly optimize the NN high-$Q$ resonances for an $N$th-order filter. However, for a random initial structure, it is not clear which $N$ of the initial (likely low-$Q$) eigenvalues are to be optimized. Similarly, it is not always easy to guess a reasonable initial structure with NN high-$Q$ resonances in the vicinity of the desired targets (in [6], we used intuition from circuit theory to guess a reasonable topology and its initial geometrical parameters). Furthermore, even if NN high-$Q$ poles are identified, as the optimization progresses and they shift, one needs to track them smoothly (differentiably) and this is not always practical, especially when two poles cross in the complex plane. Finally, compared to scattering solvers, eigensolvers tend to be slow, since they rely on solving a sequence of many linear systems [74].
@@ II Quasi-normal mode theory and inverse-design criteria
The QNMT developed in [5] considers a multiport scattering system, whose outgoing waves are related to the incoming excitations at frequency $\omega$ by a scattering matrix $S(\omega)$, and provides a simple analytical model for calculating $S(\omega)$ with very high accuracy using only knowledge about the system resonant modes. The $N$ system high-$Q$ resonances (QNMs) have complex frequencies $\omega_{n}$, which appear as poles of $S$, and complex coupling coefficient to each port $p$ equal to $D_{pn}$, which can be computed as an overlap surface integral [57] between the $n$-QNM field and the $p$-port mode at the boundary of the scatterer, as explained in detail in [5]. Conveniently, the ratios $D_{pn}/D_{qn}$ do not depend on the normalization of the QNMs ([5], Appendix D). QNMT factors these NN high-$Q$ resonances into a pole-expansion sum, indicated as $\bar{S}(\omega)$ below. Additionally, low-$Q$ resonances (with poles $\omega_{n}^{C}$ and couplings $\sigma_{n}^{C}$) can be factored into an effective-background matrix $C(\omega)$, which varies slowly with frequency (approximately constant in most practical design cases). For the case of a 2-port system, the ratios $\sigma_{n}=D_{2n}/D_{1n}$ were defined and a schematic is shown in Fig. 1. The key result from [5] is that, for a lossless system, by enforcing energy conservation (unitary $S$) and maximizing reciprocity (fine-tuning $\sigma_{n}$ to also best symmetrize $S$ — exactly symmetric when $C$ constant), the $2\times 2$ scattering matrix $S(\omega)$ has a simple, approximate, but very accurate expression that only depends on $\{\omega_{n},\sigma_{n}\}$ and $C$:
@@
(Here, $\bar{S}$ is defined as the indicated sum; note that a bar $\bar{\square}$ does not denote complex conjugation of $\square$, for which we use an asterisk $\square^{*}$.) The background $C$ is itself unitary symmetric and could conceptually be computed from the remaining (presumably low-$Q$) system poles: $C(\omega)=-\bar{S}_{\{\omega_{n}^{C},\sigma_{n}^{C}\}}(\omega)$. When no separate background is needed (all modes are considered as high-$Q$), then C=IC=-I. Moderate losses can also be incorporated perturbatively [5] by adjusting the poles only in the denominator of the $\bar{S}$ expansion in Eq. (1).
@@
[…]
For particular spectral responses, however, that correspond to standard $N$th-order filters (Butterworth, Chebyshev, inverse Chebyshev, and elliptic), described by rational functions of $i\omega$ with real coefficients [20], [6] has shown that the target QNMT parameters conveniently take specific analytical values: (i) The NN high-$Q$ complex poles $\tilde{\omega}_{n}$ are immediately available from the denominator of the desired-filter rational function [20], (ii) their coupling ratios $\tilde{\sigma}_{n}$ must have unity magnitude and alternate signs,
@@
[…]
@@
[…]
@@ III.2 Enforcing background matrix
Another optional tool at the designer’s disposal to reduce the number of undesired high-$Q$ modes is to additionally limit the total amount of material in the structure, since the density of modes increases with the material volume. If $\rho_{\text{p}}(\mathbf{r})$ denotes the (dimensionless) TopOpt “density” parameter (ranging between 0 and 1) at each location $\mathbf{r}$ in the design region $V$, the total material volume is $V_{\rho}=\int_{V}\rho_{\text{p}}(\mathbf{r})dV$. Then, this additional limit can be written as $V_{\rho}\leq V_{\rho,\text{max}}$ or, by using a slack variable $z$ to convert the inequality constraint to an equality constraint plus bounds for the variable,
@@ IV.1 Physical model of metasurface scattering
$\nabla\cdot\frac{1}{\varepsilon(x,y)}\nabla-k_{o}^{2}\mu]H_{z}=J$ (8)
@@
Both constraints of Eqs. (5) and (6) require computing the amplitudes of the outgoing waves $s_{-p}$ ($p=1,2$), by evaluating the overlap surface integral [57] between the field and the $p$-port mode at the closest cross section of the port to the scattering filter (see discussion in [5] for this proper choice of cross sections). For our metasurface structure with normally propagating plane waves, they can be calculated as $s_{-p}=\int_{0}^{a/2}H_{z}^{\text{scat}}(x_{p},y)dy$, where $x_{1},x_{2}$ are the $x$-limits of the design region and $H_{z}^{\text{scat}}(x_{p},y)$ is the scattered magnetic field, namely the total field minus the incident field. Subsequently, in taking the derivative of $s_{-p}$ (hence our objectives) with respect to the unknown design parameters, the adjoint method transforms the derivative of $s_{-p}$ with respect to $H_{z}$ into a line-source term in the adjoint equations [62].
@@ IV.2 Density-based topology optimization
Figure 3: Designed photonic 3rd-order elliptic bandpass filter: (a) The structure is shown on the top left, with the design region boxed inside the dashed orange line. Its transmission spectrum $|S_{21}(f)|^{2}$ is plotted as the orange curve, with an enlargement of the passband on the top right. Its background transmission $|C_{21}(f)|^{2}$ is below 70-70dB (and hence omitted from the plot). Therefore, the response of our design matches perfectly the target standard filter (black curve). The traditional naive-optimization method performs very poorly in comparison (green curve). (b) Target and designed poles and their coupling ratios $\{\omega_{n},\sigma_{n}\}$.
@@
$r_{\text{f}}^{2}\nabla^{2}\rho_{\text{f}}+\rho_{\text{f}}=\rho,\quad\text{with}\;\vec{n}\cdot\nabla\rho_{\text{f}}=0\;\text{on}\;{\partial\Omega}$ (9)
@@
The specifications of all transmission filters designed below are: bandpass type with $1\%$ pass bandwidth around $\lambda=1$, 0.250.25dB transmission ripple allowed in the passband, and at least 2525dB transmission suppression in the stopband. The two dielectric materials considered for the structure are $n_{\text{low}}=1$ (air) and $n_{\text{high}}=3.4$ (silicon). The incidence and transmission air regions are $0.5\lambda$ thick, which is also the thickness of the PMLs, with absorption large enough to reduce reflections to $10^{-12}$ at the left and right boundaries. To make the optimization of this bandpass filter easier, the initial structure was always chosen to be a quarter-wave Bragg stack at $\lambda=1$, namely alternating layers of $0.25$-thick air and $0.25/n_{\text{high}}$-thick silicon, so that the background transmission is close to its desired small value. To obtain reasonable computational accuracy, we chose a mesh density of about 6000 triangular mesh elements per $\lambda^{2}$ (so the average side of the triangular elements is roughly $0.02\lambda$, or $0.07\lambda/n_{\text{high}}$ in the material).
@@
As a first example, we designed a third-order ($N=3$) elliptic filter. For an odd-order filter, $\tilde{t}=0$, so the $C$ constraint in Eqs. (4) and (6) becomes $C_{21}(\omega_{m})\rightarrow 0$. The design region is chosen as $3\lambda$ long and $0.25\lambda$ high ($a=0.5\lambda$). The background transmission was constrained at $M=21$ frequency points, equally spaced within $0.85,1.15]/\lambda$. After trying multiple values for $\alpha$ and $\varphi$, we obtained the best optimization result for $\alpha=0.02$ and $\varphi=\pi/2$, namely the coupling ratios are $\sigma_{n}=\{i,-i,i\}$ from Eq. (2). The relative errors of the actual poles and coupling ratios, compared with the targets, are within $10^{-5}$. Moreover, $C_{21}$ is less than 70-70dB, effectively zero, for a free-spectral range of $30\%$. Therefore, as shown in Fig. 3, we had an exceptionally good match between the actual transmission and the target spectrum.
@@
Figure 4: Designed photonic 4th-order elliptic bandpass filter: (a) The structure is shown on the top left. Its transmission spectrum $|S_{21}(f)|^{2}$ is plotted in orange, with an enlargement of the passband on the inset. Its background transmission $|C_{21}(f)|^{2}$ is steady around the specified 25-25dB. Again, our design exhibits perfect match to the target standard filter (black curve). (b) Poles of the designed structure: the high-Q poles and their coupling ratios $\{\omega_{n},\sigma_{n}\}$ shown in the enlarged inset match their targets, while the low-Q modes collectively account for the desired background transmission.
@@
As a second example, we designed a fourth-order ($N=4$) elliptic filter, in which (even-$N$) case the 25-25dB stopband limit sets the background transmission target to the non-trivial value $\tilde{t}=10^{-25/20}=0.0562$. We chose $\varphi=-\pi/2$, so the coupling ratios are $\sigma_{n}=\{-i,i,-i,i\}$ from Eq. (2) and the target background reflection coefficient is $\tilde{r}=0.9984+\text{i}0$ from Eq. (3). As before, the design region is made $3\lambda$ long and $0.25\lambda$ high ($a=0.5\lambda$). For the background constraint, we used $M=92$ frequencies equally spaced within $0.8,1.2]/\lambda$, with $\alpha=1/150/\sqrt{M}$. After optimization, again the actual poles and coupling ratios have reached their targets to within $10^{-5}$, while $|C_{21}|^{2}$ is almost exactly at 25-25dB for a $40\%$ FSR. As shown in Fig. 4, again we have an amazing match of the actual transmission to the target spectrum.
@@ IV.4 Comparison to naive direct design
In order to confirm the superiority of our proposed optimization method, we also attempt the conventional method of directly optimizing (for the same design region and mesh) the transmission spectrum. In particular, for the third-order elliptic filter of Fig. 3, we set the optimization objective as $\min_{\rho}\int(|S_{21}^{\text{target}}|^{2}-|S_{21}^{\text{actual}}|^{2})^{2}d\omega$, minimizing the least-squares difference between the target SF transmission spectrum and the actual transmission spectrum. The integration was numerically evaluated with Gauss quadrature at 95 points within the frequency range $0.9,1.1]/\lambda$. (The increase in points and decrease in range were in recognition of the difficulty of this approach.) The optimization was performed with the CCSA algorithm [73] from NLopt [39].
@@ V Few-parameter optimization
Figure 5: Designed 1D-layered photonic 3rd-order Chebyshev bandpass filter: The dielectric-layer stack is shown on the left, where black is Si and grey is SiO2, so light is normally incident from the top air and transmits into the bottom silica substrate; the $28$ layer thicknesses (from top to bottom, alternating SiO2–Si) are $d_{k}/\lambda=(0.3528,$ $0.07358,$ $0.1787,$ $0.07361,$ $0.3449,$ $0.08524,$ $0.1795,$ $0.07385,$ $0.1793,$ $0.07383,$ $0.1794,$ $0.07391,$ $0.1804,$ $0.03658,$ $0.04277,$ $0.07453,$ $0.1794,$ $0.07382,$ $0.1792,$ $0.07380,$ $0.1793,$ $0.07385,$ $0.1797,$ $0.1212,$ $0.2876,$ $0.07501,$ $0.1854,$ $0.2154)$. The transmission spectrum $|S_{21}(f)|^{2}$ (orange) matches the target filter (black) for a wide frequency range around the passband (enlarged in the inset). The background transmission $|C_{21}(f)|^{2}$ is below 58-58dB for the entire range.
As an example, we considered such a stack of dielectric films, alternating between silicon (Si with $n_{\text{high}}=3.4$) and silicon dioxide (SiO2 with $n_{\text{low}}=1.4$), and sitting on a SiO2 substrate. We designed a third-order ($N=3$) Chebyshev filter (again bandpass with $1\%$ bandwidth and 0.250.25dB passband ripple) for plane-wave transmission from the air region above the stack into the substrate (assumed to be semi-infinite). We chose $\varphi=0$, namely the coupling ratios are $\sigma_{n}=\{1,-1,1\}$ from Eq. (2). [Note that this will force the tails of the QNMs to be (anti-)symmetric, although the structure is asymmetric.] We started with a 29-layer quarter-wave Bragg stack, namely $d_{k}=0.25\lambda/n_{k}$ for k=129k=1...29, where $n_{k}$ is the refractive index of each layer, so $d_{k=\text{odd}}=0.0735$ for Si and $d_{k=\text{even}}=0.179$ for SiO2. We bounded the individual layer thicknesses to $0.75/n_{k}$ and the total Si thickness of Eq. (13) to $d_{\rho,\text{max}}=1.5N\lambda/n_{\text{high}}$ (with multiplier $\gamma=10$), while the total thickness of Eq. (12) was left unconstrained. The initial Bragg stack already gives very low transmission, matching our desired background, and it turned out that we did not need to impose any background constraints ($M=0$).
The optimization led to a structure where the top Si layer was less than $0.01/\lambda$ thick. Therefore, we removed this layer and repeated the optimization with the remaining 28 thicknesses. The resulting transmission spectrum is shown in Fig. 5. Once again, an excellent match between the target and actual poles and passband transmission is observed. $C_{21}$ is less than 53-53dB, sufficiently close to zero, for a free-spectral range of over $40\%$.